Can a straight line pass through (2,4) , (-6,3), and (-3,1)?
Found 2 solutions by Mathtut, scott8148:
Answer by Mathtut(3670)

take any two point and write an equation find slope first....I will use the first two points change in y /change in x is the slope. 3-4/-6-2=1/8 is the slope. now take one of the points(of the two we just used to find the slope), and plug it into point slope formula(I am using the first point)
:
y-4=(1/8)(x-2)
:
we dont need to re write this we just take the third point (-3,1) and check and see if it works in the equation ....if it does it is on the same line...if not then a straight line doesn not pass thru these 3 points.
:
1-4=(1/8)(-3-2)
:
-3=1/8(-5)
:
-3=-5/8.......which is not a true statement therefore they do not all lie on the same line.

Answer by scott8148(6628)

no

arranging the x-coordinates in order __ -3 is between 2 and -6

however, looking at the corresponding y-coordinates __ 1 is NOT between 4 and 3

straight line not possible